OpenStudy (anonymous):

Could someone help me on this question, Calculate the limit algebraically: lim x-> infinity (sqroot(x^2+x) - x)

OpenStudy (anonymous):

multiply by \[\frac{\sqrt{x^2+x}+x}{\sqrt{x^2+x}+x}\]

OpenStudy (anonymous):

clear after that?

OpenStudy (anonymous):

Yeah thanks, but what would change if it was asking for limit-> -infinity?

OpenStudy (anonymous):

you would recognize that the form would not be \[\infty-\infty\] but rather \[\infty +\infty\] and the last for is clearly \[\infty\]

OpenStudy (anonymous):

*form

OpenStudy (anonymous):

Oh i get it thanks a lot, and once I simplify it would I replace x as 0 to solve?

OpenStudy (anonymous):

no you would take the limit as x goes to infinity, not zero. you will end up with a expression with a numerator and denominator, so it should be clear

OpenStudy (anonymous):

i can write it if you like, but try it

OpenStudy (anonymous):

ah ok, so I got \[x^2 / \sqrt{x^2+x} +x\] and then what would be the next step? lol sorry I'm kinda slow today :P

OpenStudy (anonymous):

oh no that is not it!

OpenStudy (anonymous):

sorry lets do the algebra

OpenStudy (anonymous):

answer should be 1/2 if my answer sheet is right :)

OpenStudy (anonymous):

yup - 1/2 is correct

OpenStudy (anonymous):

\[(\sqrt{x^2+x}-x)\times \frac{\sqrt{x^2+x}+x}{\sqrt{x^2+x}+x}\]

OpenStudy (anonymous):

we will get 1/2 eventually first we get \[\frac{x}{\sqrt{x^2+x}+x}\]

OpenStudy (anonymous):

you had an x^2 in the numerator, but it should be x

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

check the algebra, you get x^2 +x - x^2 in the numerator, leaves only x

OpenStudy (anonymous):

oh that makes sense so far

OpenStudy (anonymous):

now we use our eyeballs and think, "numerator is a poly of degree 1, and denominator is not a polynomial, but \[\sqrt{x^2}\] looks like x so the denominator behaves like \[2x\] as x goes to infinity \[\frac{x}{2x}=\frac{1}{2}\] done

OpenStudy (anonymous):

if you want to do a bunch of rather silly algebra you can multiply top and bottom by \[\frac{1}{x}\]being careful to make it look like dividing by x^2 under the radical. it is a pain and not worth writing, but if you have to turn it in it might make your teacher happy

OpenStudy (anonymous):

Oh okay, though in the numerator shouldn't it be \[\sqrt{x^2 +x}\] instead of just sqroot(x^2)?

OpenStudy (anonymous):

\[\frac{\frac{x}{x}}{\sqrt{\frac{x^2}{x^2}+\frac{x}{x^2}+\frac{x}{x}}}\]

OpenStudy (anonymous):

\[=\frac{1}{\sqrt{1+\frac{1}{x}}+1}\]

OpenStudy (anonymous):

take the limit as x goes to infinity, get \[\frac{1}{2}\] yes you are right, it is \[\sqrt{x^2+x}\] in the denominator, that is why i said i was using the eyeball method. ignoring the "x" term

OpenStudy (anonymous):

ohhh okay then

OpenStudy (anonymous):

in other words thinking "as x goes to infinity the square root of x squared plus x looks like the square root of x squared (because the x is unimportant compared to the x squared) and the square root of x squared looks like x

OpenStudy (anonymous):

OH, I get it now, thanks a lot for your help and time, I appreciate it :D

OpenStudy (anonymous):

yw